OpenStudy (anonymous):

f(x) = a(x+1)(x+7-i)(x+7+i) factored?

OpenStudy (anonymous):

this is already factored, do you want to multiply it out?

OpenStudy (anonymous):

yes :) sorry for the wrong wording

OpenStudy (anonymous):

F(x)=a(x+1)[(x+7)^2-i^2] =a(x+1)[x^2+49 -14x+1] =a(x+1)[x^2-14x+50] =a{x^3-14x^2+50x+x^2-14x+50] =a{x^3-13x^2+36x+50}

OpenStudy (anonymous):

thank you i was struggling with that, i am not good at that for some reason

OpenStudy (anonymous):

Not a problem.